9. A researcher is interested in whether child aggressive behavior changes after they see another child who has something that they want (but do not have). She devises an experiment where children come into her lab and are given worn and broken toys to play with. After a short time, another child is brought in the room and is given new, shiny toys to play with. The researcher wants to know whether there are changes in aggressive behavior among the children with the broken, worn toys after they see the other children playing with new, shiny toys. Therefore she measures the number of aggressive behaviors among the original children given the worn and broken toys before and after they saw the other children playing with new, shiny toys. Her findings are summarized in the table below. Subject Before After 18 16 10 17 4 19 10 12 14 a. Identify the independent and dependent variables. What are the levels of the IV? b. Is this experiment a paired-samples or independent-samples design? c. State null and alternative hypotheses. d. Conduct a statistical test of the hypothesis. f. Interpret your results. Be sure to make a decision about whether to reject or retain the null hypothesis.
Answer #1

a) Independent variable: Number of different toys

Dependent variable: Number of aggressive behaviors

b) Experiments is paired samples because data is given before and after type.

c) The null and alternative hypothesis is

H0:\mu_{d}=0

H1:\mu_{d}\neq 0

Level of significance = 0.05

d)

Subject Before After Difference(d)
A 16 18 -2
B 10 11 -1
C 17 19 -2
D 4 6 -2
E 9 10 -1
F 12 14 -2

Sample size = n = 6

The sample mean of difference = \bar{d} = - 1.6667

Sample standard deviation of difference = s_{d} = 0.5164

Test statistic is

t=\frac{\bar{d}}{s_{d}/\sqrt{n}}

t=\frac{-1.6667}{0.5164/\sqrt{6}}=-7.91

Degrees of freedom = n - 1 = 6 - 1 = 5

Critical value = 2.571 ( Using t table)

Test statistic | t | > critical value we reject null hypothesis.

f) Conclusion: There are changes in aggressive behavior among children with broken worm toys after they see other children playing with new, shiny toys.